# Strategic Informed Trading and the Value of Private Information

### v1

@@ 3. Welfare Analysis
Two quite surprising consequences stem from the above proposition. First, that the insider’s welfare might be higher in the PT equilibrium is directly attributable to the presence of the private signal, as when there is no private signal, internalization of price impact is always beneficial for the insider. Second, assuming no private signal and that both $I,U$ have the same risk tolerance, which forces $\lambda=1/2$ we have the following almost sure order of interim certainty equivalents
@@
Amazingly, not only it is better for the uninformed agent when the insider internalizes her price impact, the uninformed agent’s welfare actually exceeds that of the insider. We explain the mechanism behind these predictions in the next section.
4. Equilibria Structure
In this section, we analyze and compare the equilibrium quantities (allocation and prices) in order to infer the economic intuition behind the models predictions on the effects of price-impact internalization (“internalization”) and insider’s private signal (“private information”). As will be shown, internalization and private information may have competing affects on the insider’s demand, and hence equilibrium quantities. Broadly, internalization has the effect of dampening the position size, while private information may increase the insider’s equilibrium allocation of risk.
@@ 4. Equilibria Structure
Private information effects when price taking
@@
(37) $\widehat{\psi}_{I}(g,z)-\widehat{\psi}_{ns,I}(z)=\frac{(1-\lambda)(p_{I}-p_{U})}{1+\lambda p_{I}+(1-\lambda)p_{U}}(g-p_{0})-\frac{(1-\lambda)(p_{I}(\lambda p_{I}+(1-\lambda)p_{U})+p_{U})}{p_{I}(1+\lambda p_{I}+(1-\lambda)p_{U})}\frac{z}{\alpha_{I}}.$
@@
Another interesting effect concerns the outstanding supply, $\Pi=(\alpha_{I}+\alpha_{U})\widehat{\Pi}=-(\alpha_{I}+\alpha_{U})p_{0}$. At the ex-ante level, the expected demand change is
@@
so that on average the insider’s position increases due to the private signal. Intuitively, positive supply means that the insider and the uniformed trader are expected to buy the tradeable asset at the equilibrium and for the insider the presence of private signal means that her estimated variance of the tradeable asset (i.e. its risk) is lower. This increases her demand, implying that ex-ante she is more confident to hold higher part of asset supply.
@@
$p(h(g,z))-p_{ns}(z)=\frac{\lambda p_{I}+(1-\lambda)p_{U}}{1+\lambda p_{I}+(1-\lambda)p_{U}}(g-p_{0})+\frac{\lambda[(1-\lambda)(p_{I}-p_{U})-p_{I}(consbp_{I}+(1-\lambda)p_{U})]}{p_{I}(1+\lambda p_{I}+(1-\lambda)p_{U})}\frac{z}{\alpha_{I}}.$
This shows the equilibrium price is increasing in the signal (the factor in front of $z/\alpha_{I}$ can be positive or negative). At the ex-ante level
@@
Thus, the presence of the signal increases both the insider’s expected demand and expected equilibrium price. Additionally, the equilibrium clearing condition (4) implies the private signal has the opposite effect on the uniformed trader’s demand [^13] The uninformed agent sees $Z_{N}$ (through the price) in the no-signal equilibrium, but does not see $Z_{N}$ in the PT equilibrium. Therefore, we are not saying the uninformed agent sees $Z_{N}$ in the PT equilibrium and then adjusts her position accordingly. Rather, we are saying the effect of noise trading in the PT equilibrium is to increase the trade size of the uninformed agent over the no-signal equilibrium.. Summarizing,
The presence of private information in the PT equilibria is expected to increase the insider’s demand and the price and to decrease the uniformed trader’s demand.
@@
As in the PT equilibrium, market clearing implies the effects of the signal on the uniformed trader’s demand are in thee opposite direction of the insider. Indeed, direct calculation shows
@@
which means that in contrast to the insider, the uniformed trader’s demand is expected to decease due to price impact. Thus, we may conclude
For both the PT and PI equilibria, the presence of a private signal is expected to increase the insider’s demand and price (albeit with a lower change in the PI equilibrium) and decrease the uniformed trader’s demand.
@@
The above implies that in the PI equilibrium, the insider has a lower demand when compared to the PT equilibrium and obtains a better price (price-impact increases the per-unit price when insider sells and decreases it when she buys). Indeed, positive $z$ means that both insider and uniformed trader sell at equilibrium. In the view of Remark 2.2, positive $z$ makes the insider reveal a higher public signal. This increases the demand of the uniformed trader (covers less of noise demand) and hence the interim effect is an increase in the equilibrium price. Lastly, the uniformed trader’s equilibrium allocations imply the relative trade size
@@
Therefore, when $z<0$ both the insider and uninformed trader buy the risky asset in each equilibria. However, in the PI equilibrium, the insider reduces her position while the uninformed trader increases it. Conversely, when $z>0$ traders sell the risky asset, with the uniformed trader selling more. In particular, the uninformed increases his volume at a better price-per-unit, due to price impact. In other words, when insider buys the asset, internalization reduces her demand which in turn decreases the price and makes the uniformed trader buys more. Summarizing,
Internalization with no signal results in a lower (resp. higher) equilibrium position for the insider (resp. uniformed trader) at a better price.
@@
where the second equality follows using (29) and simplifying. The following shows the right side above may be either positive or negative. In short, the effect of price impact prevails over the one of asymmetric information when the insider is sufficiently risk tolerant resulting is a higher expected order.
@@
Proposition 2.4 implies the right side above is negative, which means that internalization is expected to decrease (resp. increase) the price when insider is expected to buy (resp. sell). In other words, the expected change of the price benefits both traders (as the uniformed trader remains at the same side of trade). Connecting this fact to the expected demand changes, we may conclude that
@@
Due to internalization of price impact, a sufficiently low (resp. high) risk tolerant insider is expected to buy less (resp. more) units at a better price, while uniformed trader buys more (resp. less).
Intuition and model’s predictions
@@
We conclude by providing economic intuition about the predictions induced by the model. First, we have seen that asymmetric information is ex-ante expected to increase the insider’s demand for the tradeable asset. Intuitively, the private signal reduces the asset’s risk (measured by variance) for the insider, which makes her willing to hold a higher allocation. Without strategic trading, higher insider demand is expected to increase the price. Private information tends to increase the insider’s demand even under price impact. The main ex-ante expected difference is the lower increase of the price, as internalization of price impact affects the price in favor of the insider. When the insider trades strategically, she uses her private signal to affect the equilibrium. In fact, it is exactly when the insider internalizes price impact and uninformed agents are price-takers that insider’s welfare is monotonically increasing with respect to the signal precision. In other words, insider’s strategic trading increases the value of her private signal making the acquisition of better signal reasonable.
@@
Internalization of price impact keeps the traders at the same side of trade. As we have seen, the insider hides part of her private signal, which is expected to reduce her demand and hence the equilibrium price. In other words, price impact has the opposite expected effect than the private signal. Provided that insider has long position at equilibrium, the price always decreases due to price impact, under both price impact and private signal. However, the insider’s demand is lower under the presence of signal when she is also sufficiently risk averse. This is intuitive in the sense that a sufficiently risk averse trader wants to undertake less risk. If in addition to high insider’s risk aversion inequality (36) holds, the price impact equilibrium lowers the insider’s welfare. Note that (36) implies the uniformed traders are highly risk tolerant. This reduces the insider’s allocation even further, since the uniformed traders’ demand is higher and hence at market-clearing the insider’s share is lower. Such lower insider’s demand may lead to lower ex-ante expected utility gains, as a signal of good quality (consistent with (36)) means the insider feels more confident to hold a higher position at equilibrium. However, under a large deviation of risk aversions, internalization of price impact prevails, demand is reduced and utility gains are lower. Note that when traders have the same risk aversion, the effect of internalization is always beneficial for the insider.
@@
Interestingly enough, in the absence of a private signal, internalization of price impact induces higher utility gains for the insider. This is because under symmetric information, the insider does not have motive to increase her demand due to a lower asset risk, and hence the effect of price impact, i.e. buying lower quantity at a lower price, increases the expected welfare. We conclude that it is the presence of asymmetric information and the deviation on risk aversions that potentially make the price-impact equilibrium disadvantageous for the insider.
@@
We should also emphasize that as long as the insider internalizes her price impact, equilibrium prices cannot be driven to the corresponding PT equilibrium. This is because the uniformed trader, although he is a price-taker, realizes the insider internalizes her price impact. This makes him reduce his perceived public signal precision, and hence alters his demand function to a less elastic one. Under the Pareto initial allocation, he remains at the same side of trade with the insider and the lower equilibrium prices caused by the internalization imply higher demand for the uniformed trader, who (although price-taker) is benefited by price impact. In other words, price impact decreases the price when traders buy the asset, and at equilibrium the uniformed trader satisfies his optimal demand but at a discount. This is the reason why price impact ex-ante benefits the uniformed trader with and without asymmetric information.
@@ Appendix A The Equilibrium: General Case
We extend the model of Section 1 to $d$ risky assets with general values for the model inputs, outstanding supply, and agent initial endowments. The assets have terminal payoff $X\sim N(\mu_{X},P_{X}^{-1})$, written
@@
where $\mathcal{E}_{X}\sim N(0,1_{d})$. The outstanding supply is $\Pi\in{\mathbb{R}}^{d}$. The riskless asset price is still normalized to $1$. The private signal $G$ takes the form
@@
where $\mathcal{E}_{n}\sim N(0,1_{d})$ is independent of both $X$ and $\mathcal{E}_{I}$. The matrices $P_{X},P_{I},P_{N}$ lie in $\mathbb{S}^{d}_{++}$, the set of $d\times d$ strictly positive definite symmetric matrices, and $\mu_{X}\in\mathbb{R}^{d}$. Lastly, the insider $I$ and uninformed agent $U$ are endowed with share positions $\left\{\pi_{i,0}\right\}$ consistent with equilibrium in that (with $\pi_{i,0}=\alpha_{i}\psi_{i,0}$) $\Pi=\alpha_{I}\psi_{I,0}+\alpha_{U}\psi_{U,0}$.
@@
For a given strategy $\pi_{i}$, if the time 00 price vector is $p$ then agent $i$ has risk aversion adjusted terminal wealth (see (5))
(46) $\mathcal{W}^{\psi_{i}}=(\psi_{i,0})^{\prime}p+\psi_{i}^{\prime}(X-p);\qquad i\in\left\{I,U\right\}.$
where the symbol denotes transposition. The insider perceives linear price impact so that if she changes her position from $\pi_{I,0}=\alpha_{I}\widehat{\Pi}$ to $\pi_{I}=\alpha_{I}\psi_{I}$, the price will be as in (6) (for a to-be-determined vector $V$ and matrix $M$). The set of acceptable strategies for the insider ${\mathcal{A}}_{I}$ is the same as in (7) and the insider’s optimal investment problem is
@@
(47) $\begin{split}&\inf_{\psi\in{\mathcal{A}}_{I}}\mathbb{E}\left[e^{-\left(\psi_{I,0}^{\prime}p_{\iota}\left(\psi,Z_{N}\right)+\psi^{\prime}\left(X-p_{\iota}\left(\psi,Z_{N}\right)\right)\right)}\big{|}\sigma(G,Z_{N})\right].\end{split}$
@@
Write $P_{X|G}$ as the precision of $X$ given the insider signal $G$, and similarly to (9) express $M$ in terms of $P_{X|G}$ and a to-be-determined matrix ${\mathcal{Y}}$.
(48) $\begin{split}P_{X|G}:=P_{I}+P_{X};\qquad M=M({\mathcal{Y}})=P_{X|G}^{-1/2}\ {\mathcal{Y}}\ P_{X|G}^{-1/2}.\end{split}$
@@
We also write $V=V({\mathcal{Y}})$ to stress the dependence, and expect ${\mathcal{Y}}+{\mathcal{Y}}^{\prime}\in\mathbb{S}^{d}_{++}$. Analogously to Lemma 1.1 we obtain a unique solution to the insider’s optimal investment problem
@@
Let $g,z\in\mathbb{R}$. On the set $\left\{G=g,Z_{N}=z\right\}$, the unique optimizer of (47) enforces
(49) $\begin{split}\widehat{\psi}_{I,\iota}(g,z)-\psi_{I,0}+\frac{z}{\alpha_{I}}&={\mathcal{M}}\left(g+\Lambda_{\iota}z+\mathcal{V}\right),\end{split}$

### v2

@@ 3. Indirect Utility
Two quite surprising consequences stem from the above proposition. First, that any utility gain the insider obtains in PT equilibrium over the PI equilibrium is attributable solely to the presence of the private signal, as when there is no private signal, internalization of price impact is always beneficial. Second, assuming no private signal and that $I,U$ have the same risk tolerance, (which implies $\lambda=1/2$) we have the almost sure order of interim certainty equivalents
@@
Amazingly, not only it is better for the uninformed trader when the insider internalizes her price impact, the uninformed trader’s utility actually exceeds that of the insider. We explain the mechanism behind these predictions in the next section.
4. Equilibrium Structure
In this section, we analyze and compare the equilibrium quantities (allocation and prices) in order to infer the economic intuition behind the model’s predictions on the effects of price-impact internalization (“internalization”) and insider’s private signal (“private information”). As will be shown, internalization and private information may have competing affects on the insider’s demand, and hence equilibrium quantities. Broadly, internalization has the effect of dampening the position size, while private information may increase the insider’s equilibrium allocation of risk.
@@ 4. Equilibrium Structure
(30) $\widehat{\psi}_{I}(g,z)-\widehat{\psi}_{ns,I}(z)=\frac{(1-\lambda)(p_{I}-p_{U})}{1+\lambda p_{I}+(1-\lambda)p_{U}}\left(g-p_{0}-\frac{p_{I}(\lambda p_{I}+(1-\lambda)p_{U})+p_{U}}{p_{I}(p_{I}-p_{U})}\frac{z}{\alpha_{I}}\right).$
@@
Another interesting effect concerns the outstanding supply, $\Pi=(\alpha_{I}+\alpha_{U})\widehat{\Pi}=-(\alpha_{I}+\alpha_{U})p_{0}$. As both $G,Z_{N}$ have mean 00, at the ex-ante level, the expected demand change is
@@
so that on average the insider’s position increases due to the private signal when the supply is positive. Intuitively, a positive supply means agents are expected to buy the asset in equilibrium, and for the insider the private signal lowers the variance of the asset payoff (i.e. its risk) is lower. This increases her demand, implying ex-ante that she is willing to hold a larger position.
@@
$\begin{split}p(h(g,z))-p_{ns}(z)&=\frac{\lambda p_{I}+(1-\lambda)p_{U}}{1+\lambda p_{I}+(1-\lambda)p_{U}}\bigg{(}g-p_{0}+\lambda\bigg{(}\frac{(1-\lambda)(p_{I}-p_{U})}{p_{I}(\lambda p_{I}+(1-\lambda)p_{U})}-1\bigg{)}\frac{z}{\alpha_{I}}\bigg{)}.\end{split}$
This shows the equilibrium price is increasing in the insider signal, but is ambiguous in the noise as the factor in front of $z/\alpha_{I}$ can be positive or negative. At the ex-ante level
@@
Thus, the presence of the signal increases both the insider’s expected demand and expected equilibrium price. Additionally, the equilibrium clearing condition (3) implies the private signal has the opposite effect on the uninformed trader’s demand [^12] The uninformed trader sees $Z_{N}$ (through the price) in the no-signal equilibrium, but does not see $Z_{N}$ in the PT equilibrium. Therefore, we are not saying the uninformed trader sees $Z_{N}$ in the PT equilibrium and then adjusts her position accordingly. Rather, we are saying the effect of noise trading in the PT equilibrium is to increase the trade size of the uninformed trader over the no-signal equilibrium.. Summarizing,
The presence of private information in the PT equilibria is expected to increase the insider’s demand and the price and to decrease the uninformed trader’s demand.
@@
As in the PT equilibrium, market clearing implies the effects of the signal on the uninformed trader’s demand are in thee opposite direction of the insider. Indeed, direct calculation shows
@@
which means that in contrast to the insider, the uninformed trader’s demand is expected to decrease due to price impact. Thus, we may conclude
For both the PT and PI equilibria, the presence of a private signal is expected to increase the insider’s demand and price (albeit with a lower change in the PI equilibrium) and decrease the uninformed trader’s demand.
@@
The above implies that in the PI equilibrium, the insider has a lower demand when compared to the PT equilibrium and obtains a better price (price-impact increases the per-unit price when insider sells and decreases it when she buys). Indeed, positive $z$ means that both insider and uninformed trader sell at equilibrium and in fact positive $z$ makes the insider reveal a higher public signal [^13] From (23) we get that price impact changes the public signal in the direction of the noise traders’ order. Indeed, as $\widehat{y}>0$, we have $(h_{\iota}(g,z)-h(g,z))z\geq 0$ for all $(g,z)$. Hence, the strategically revealed signal by the insider is higher if and only if there is positive demand from the noise traders.. This increases the demand of the uninformed trader (covers less of noise demand) and hence the interim effect is an increase in the equilibrium price. Lastly, the uninformed trader’s equilibrium allocations imply the relative trade size
@@
Therefore, when $z<0$ both the insider and uninformed trader buy the risky asset in each equilibria. However, in the PI equilibrium, the insider reduces her position while the uninformed trader increases it. Conversely, when $z>0$ traders sell the risky asset, with the uninformed trader selling more. In particular, the uninformed increases his volume at a better price-per-unit, due to price impact. In other words, when insider buys the asset, internalization reduces her demand which in turn decreases the price and makes the uninformed trader buys more. Summarizing,
Internalization with no signal results in a lower (resp. higher) equilibrium position for the insider (resp. uninformed trader) at a better price.
@@
where the second equality follows using (22) and simplifying. The following shows the right side above may be either positive or negative. In short, the effect of price impact prevails over the one of asymmetric information when the insider is sufficiently risk tolerant resulting in a higher expected order.
@@
Proposition 2.3 implies the right side above is negative, which means that internalization is expected to decrease (resp. increase) the price when insider is expected to buy (resp. sell). In other words, the expected change of the price benefits both traders (as the uninformed trader remains at the same side of trade). Connecting this fact to the expected demand changes, we may conclude that
@@
Due to internalization of price impact, a sufficiently low (resp. high) risk tolerant insider is expected to buy less (resp. more) units at a better price, while uninformed trader buys more (resp. less).
5. Intuition and model’s predictions
@@
We conclude by providing economic intuition about the predictions induced by the model. First, we have seen that asymmetric information is ex-ante expected to increase the insider’s demand for the tradeable asset. Intuitively, the private signal reduces the asset’s risk (measured by variance) for the insider, which makes her willing to hold a higher allocation (positive or negative). Without strategic trading, higher insider demand is expected to increase the price. Private information tends to increase the insider’s demand even under price impact. The main ex-ante expected difference is the lower increase of the price, as internalization of price impact affects the price in favor of the insider. When the insider trades strategically, she uses her private signal to affect the equilibrium. In fact, it is exactly when the insider internalizes price impact and uninformed traders are price-takers that insider’s utility is monotonically increasing with respect to the signal precision. In other words, insider’s strategic trading increases the value of her private signal, making the acquisition of a better signal reasonable.
@@
Internalization of price impact keeps the traders at the same side of trade. As we have seen, the insider hides part of her private signal, which is expected to reduce her demand and hence the equilibrium price. In other words, price impact has the opposite expected effect than the private signal. Provided that insider has long position at equilibrium, the price always decreases due to price impact, under both price impact and private signal. However, the insider’s demand is lower under the presence of signal when she is also sufficiently risk averse. This is intuitive in the sense that a sufficiently risk averse trader wants to undertake less risk. If in addition to high insider’s risk aversion inequality (29) holds, the price impact equilibrium lowers the insider’s utility. Note that (29) implies the uninformed traders are highly risk tolerant. This reduces the insider’s allocation even further, since the uninformed traders’ demand is higher and hence at market-clearing the insider’s share is lower. Such lower insider demand may lead to lower ex-ante expected utility gains, as a signal of good quality (consistent with (29)) means the insider feels more confident to hold a higher position at equilibrium. However, under a large deviation of risk aversions, internalization of price impact prevails, demand is reduced and utility gains are lower. Note that when traders have the same risk aversion, the effect of internalization is always beneficial for the insider.
@@
Interestingly enough, in the absence of a private signal, internalization of price impact induces higher utility gains for the insider. This is because under symmetric information, the insider does not have motive to increase her demand due to a lower asset risk, and hence the effect of price impact, i.e. buying lower quantity at a lower price, increases the expected utility. We conclude that it is the presence of asymmetric information and the deviation on risk aversions that potentially make the price-impact equilibrium disadvantageous for the insider.
@@
We should also emphasize that as long as the insider internalizes her price impact, equilibrium prices cannot be driven to the corresponding price-taking equilibrium. This is because the uninformed trader, although he is a price-taker, realizes the insider internalizes her price impact. This makes him reduce his perceived public signal precision, and hence alters his demand function to a less elastic one. Under the Pareto initial allocation, he remains at the same side of trade with the insider and the lower equilibrium prices caused by the internalization imply higher demand for the uninformed trader, who (although price-taker) is benefited by price impact. In other words, price impact decreases the price when traders buy the asset, and at equilibrium the uninformed trader satisfies his optimal demand but at a discount. This is the reason why price impact ex-ante benefits the uninformed trader with and without asymmetric information.
@@ Appendix A The Equilibrium: General Case
In this appendix, we extend the model of Section 1 to $d$ risky assets with general values for the payoff mean and covariance. We also precisely quantify the affine coefficients in the optimal demand functions, and derive the cubic equation (6). The assets have terminal payoff $X\sim N(\mu_{X},P_{X}^{-1})$, written
@@
where $\mathcal{E}_{X}\sim N(0,1_{d})$. The outstanding supply is $\Pi\in{\mathbb{R}}^{d}$. The riskless asset price is normalized to $1$. The private signal $G$ takes the form
@@
where $\mathcal{E}_{n}\sim N(0,1_{d})$ is independent of both $X$ and $\mathcal{E}_{I}$. The matrices $P_{X},P_{I},P_{N}$ lie in $\mathbb{S}^{d}_{++}$, the set of $d\times d$ strictly positive definite symmetric matrices, and $\mu_{X}\in\mathbb{R}^{d}$. Lastly, the insider $I$ and uninformed trader $U$ are endowed with share positions $\left\{\pi_{i,0}\right\}=\alpha_{i}\widehat{\Pi},i\in\left\{I,U\right\}$ exactly as in (1).
@@
For a given strategy $\pi_{i}$, if the time 00 price vector is $p$ then trader $i$ has risk aversion adjusted terminal wealth (see (4))
$\mathcal{W}^{\psi_{i}}=(\psi_{i,0})^{\prime}p+\psi_{i}^{\prime}(X-p)=\widehat{\Pi}^{\prime}p+\psi_{i}^{\prime}(X-p);\qquad i\in\left\{I,U\right\}.$
where the symbol denotes transposition. The insider perceives linear price impact so that if she changes her position from $\pi_{I,0}=\alpha_{I}\widehat{\Pi}$ to $\pi_{I}=\alpha_{I}\psi_{I}$, the price will be as in (5) (for a to-be-determined vector $V_{p,\iota}$ and matrix $M_{p,\iota}$). To define the set of acceptable strategies for the insider ${\mathcal{A}}_{I}$, note that $G$ and $Z_{N}$ are the only random quantities revealed at time 00, and hence every insider strategy must be known using the information generated by $G$ and $Z_{N}$. Next, recall that our assumption (as in Rochet and Vila [1994]) is that the insider, by seeing both her private signal and the resultant price, is able to deduce the noise trader demand. But, if the insider uses a strategy $\psi$ which reveals the noise trader demand $Z_{N}$ through the price, it must also be that $Z_{N}$ is known given the information generated by $G$ and $p_{\iota}(\psi,Z_{N})$. As such, we define set of acceptable trading strategies for the insider to be
Here, “$\psi\in\sigma(G,H)$” means that $\psi$ is $\sigma(G,H)$ measurable, and note that for any $\psi\in{\mathcal{A}}_{I}$, one has $\sigma(G,p_{\iota}(\psi,Z_{N}))=\sigma(G,Z_{N})$. With ${\mathcal{A}}_{I}$ property defined, the insider’s optimal investment problem is
@@
(41) $\begin{split}&\inf_{\psi\in{\mathcal{A}}_{I}}\left(\mathbb{E}\left[e^{-\left(\widehat{\Pi}^{\prime}p_{\iota}+\psi^{\prime}\left(X-p_{\iota}\right)\right)}\big{|}\sigma(G,Z_{N})\right];\quad p_{\iota}=p_{\iota}\left(\psi,Z_{N}\right)\right).\end{split}$
We next turn to the uninformed trader. To write his optimal investment problem assume there is signal $H_{\iota}$ passed to all market participants in equilibrium, and the price can be written $p_{\iota}=p_{\iota}(H_{\iota})$. Both $H_{\iota}$ and $p_{\iota}(H_{\iota})$ are precisely defined below, but for now we take them as given. As the uninformed is a price taker, his optimization problem is
(42) $\begin{split}&\inf_{\psi\in\sigma(H_{\iota})}\left(\mathbb{E}\left[e^{-\left(\widehat{\Pi}^{\prime}p_{\iota}+\psi^{\prime}\left(X-p_{\iota}\right)\right)}\big{|}\sigma(H_{\iota})\right];\quad p_{\iota}=p_{\iota}\left(H_{\iota}\right)\right).\end{split}$
Transformations
[…]
(43) $\begin{split}(I)&\qquad\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\big{|}\sigma(G,Z_{N})\right]\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\theta^{\prime}\left(X-p_{\iota}\right)}\big{|}\sigma(G,Z_{N})\right];\quad p_{\iota}=p_{\iota}\left(\theta,Z_{N}\right),\\ (U)&\qquad\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\big{|}\sigma(H_{\iota})\right]\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\theta^{\prime}\left(X-p_{\iota}\right)}\big{|}\sigma(H_{\iota})\right];\quad p_{\iota}=p_{\iota}\left(H_{\iota}\right).\end{split}$
As such, we can work under ${\mathbb{Q}}_{0}$, and using (1) the clearing condition (3) is
(44) $0=\alpha_{I}\widehat{\theta}_{I}+\alpha_{U}\widehat{\theta}_{U}+Z_{N},$
which effectively makes the outstanding supply 00. The price impact form (5) is
(45) $p_{\iota}(\theta,Z_{N})=M_{p,\iota}\left(\theta+\frac{Z_{N}}{\alpha_{I}}\right)+V_{p,\iota}.$
Next, we de-mean the terminal payoff, price and signals. To do so, extending (8), define
(46) $p_{{\mathbb{Q}}_{0}}:=\mathbb{E}^{{\mathbb{Q}}_{0}}\left[X\right]=\mu_{X}-P_{X}^{-1}\widehat{\Pi},$
as the equilibrium price absent private signals provided the initial endowments satisfy (1). With this notation, set
(47) $X_{0}:=X-p_{{\mathbb{Q}}_{0}};\qquad p_{0,\iota}:=p_{\iota}-p_{{\mathbb{Q}}_{0}};\qquad G_{0}:=G-p_{{\mathbb{Q}}_{0}};\qquad H_{\iota,0}:=H_{\iota}-p_{{\mathbb{Q}}_{0}}.$
Clearly, $\sigma(G,Z_{N})=\sigma(G_{0},Z_{N})$ and $\sigma(H_{\iota,0})=\sigma(H_{\iota})$. Also, from (38) we see $G_{0}=X_{0}+Z_{I}$ where $Z_{I}$ is a $N(0,P_{I}^{-1})$ random vector independent of $X_{0}$ under ${\mathbb{Q}}_{0}$. As such, $G_{0}$ under ${\mathbb{Q}}_{0}$ takes the same form as $G$ under ${\mathbb{P}}$, except now the payoff and insider signal have mean 00. Similarly, $Z_{N}$ is a $N(0,P_{N}^{-1})$ random vector under ${\mathbb{Q}}_{0}$ independent of both $X_{0},Z_{I}$. With this notation, from (43) we start the the insider’s problem
(48) $\begin{split}\inf_{\theta\in{\mathcal{A}}_{I}}\left(\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\theta^{\prime}\left(X_{0}-p_{0,\iota}\right)}\big{|}\sigma(G_{0},Z_{N})\right];\quad p_{0,\iota}=M_{p,\iota}\left(\theta+\frac{Z_{N}}{\alpha}\right)+V_{0,p,\iota}\right),\end{split}$
@@
where $V_{0,p,\iota}=V_{p,\iota}-p_{{\mathbb{Q}}_{0}}$. We write $P_{X|G}$ as the ${\mathbb{P}}$ precision of $X$ given $G$ (and also the ${\mathbb{Q}}_{0}$ precision of $X_{0}$ given $G_{0}$), and express the matrix $M_{p,\iota}$ in terms of $P_{X|G}$ and a to-be-determined matrix ${\mathcal{Y}}$.
(49) $\begin{split}P_{X|G}:=P_{I}+P_{X};\qquad M_{p,\iota}=M_{p,\iota}({\mathcal{Y}})=P_{X|G}^{-1/2}\ {\mathcal{Y}}\ P_{X|G}^{-1/2}.\end{split}$
@@
We also write $V_{0,p,\iota}=V_{0,p,\iota}({\mathcal{Y}})$ to stress the dependence, and expect ${\mathcal{Y}}+{\mathcal{Y}}^{\prime}\in\mathbb{S}^{d}_{++}$. With this notation we obtain
@@
Let $g,z\in\mathbb{R}$. On $\left\{G_{0}=g,Z_{N}=z\right\}$, the unique optimizer of (48) enforces
(50) $\begin{split}\widehat{\theta}_{I,\iota}(g,z)+\frac{z}{\alpha_{I}}&={\mathcal{M}}\left(g+\Lambda_{\iota}z-P_{I}^{-1}P_{X|G}V_{0,p,\iota}({\mathcal{Y}})\right),\end{split}$